OpenStudy (anonymous):

For what value of c will y= x^2 + (c/x) have a relative minimum at x=-1? a: c=-4 b: c=-2 c: c=2 d: c=4 e: none of these

OpenStudy (anonymous):

\[y=x^2+\frac{ c }{ x }\] \[\frac{ dy }{ dx }=2x-\frac{ c }{ x^2 }\] it has a relative minimum at x=-1 Hence \[\frac{ dy }{ dx }=0~at~x=-1\] \[2*-1+\frac{ c }{(- 1)^2 }=0,c=2\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw